when f(x)= x^2-3x+5 g(x)=3x-1
find f[g(-1)]


Answer by stanbon(75887)

when f(x)= x^2-3x+5 g(x)=3x-1
find f[g(-1)]
------
f[g(-1)] = f[-3-1]
= f[-4] = (-4)^2-3(-4)+5 = 16+12+5 = 33
===========================================
Cheers,
Stan H.